Quantum Holonomic Computing via Lax Type Flows on Grassmannian Manifolds and Dual

Momentum Mappings

Prykarpatsky Anatoliy Dept. of Applied Mathematics at the Academy of Mining and Metallurgy, Cracow

D. L. Blackmore

Ufuk Taneri

A general approach to holonomic quantum computing via Lax -type flows on Grassmann manifolds, based on the momentum mapping reduction and connection techniques is developed. It is shown that the associated holonomy groups can be effectively used in quantum computations of diverse practical problems.